OpenStudy (anonymous):

-3x-2<5 does any one no the answer

OpenStudy (bahrom7893):

-3x <3 x<-1

OpenStudy (anonymous):

are you sure?

OpenStudy (bahrom7893):

try it: say x is -2 -3*-2 -2 = 6 - 2 = 4<5

OpenStudy (bahrom7893):

and sure im sure

OpenStudy (bahrom7893):

sorry this is wrong...

OpenStudy (bahrom7893):

I thought it was a +2 not a minus 2

OpenStudy (anonymous):

ya i can see

OpenStudy (bahrom7893):

x<-7/3

OpenStudy (anonymous):

still wrong

OpenStudy (bahrom7893):

how?

OpenStudy (anonymous):

x>-7/3

OpenStudy (bahrom7893):

what the heck, omg i forgot, u change the direction too when u multiply by -1... this is what calculus does to u.. =( Embarassing!

OpenStudy (anonymous):

maybe there are a lot of formulae at your brain right now :)